# NACHOS: Neural Architecture Search for Hardware Constrained Early Exit Neural Networks

### v1

@@ I Introduction
Neural Architecture Search (NAS) is a new research area [17] aiming at automating the design of a neural network by exploring different architectural configurations given a neural network topology, a task to be accomplished, and a dataset used for training and validation so as to provide the optimal architecture. The majority of works in this field focus on the design of deep neural network architectures while taking into account a specific target hardware. This family of NAS solutions is called Hardware-Aware NAS [18]. An emerging direction of these NAS solutions regards the use of constraints in the NAS exploration [19, 20, 21, 22]. Specifically, the works belonging to this research direction are able to consider both functional constraints (i.e., the type of operations that can be carried out in the neural network) and technological constraints (i.e., constraints on the computational and memory demand of the designed neural network in the exploration of the neural networks). The constraints are usually handled by a genetic algorithm applying a penalization to the fitness of the architectures not satisfying the hardware constraints. Interestingly, there are very few examples of NAS for EENNs in the current literature and these works are not able to provide an effective joint design of the backbone networks and the EECs [23, 24, 21, 25]. These NAS solutions are commented in Section II.
@@
  • a novel NAS solution able to co-design the backbones and the EECs in EENNs satisfying user-defined constraints;
  • the automatic placement of the EECs in the backbone of an EENN accounting for their computational complexity;
@@
The rest of this paper is organized as follows. Section II introduces the related works regarding EENNs designed by NAS solutions. Section III provides the background. Section IV introduces the NACHOS framework developed in this study, while in Section V the experimental results aiming at evaluating the effectiveness of NACHOS are shown. Conclusions are finally drawn in Section VI. To facilitate comparisons and reproducibility, the source code of NACHOS is released to the scientific community as a public repository. [^1] https://github.com/AI-Tech-Research-Lab/NAS
@@ II Related literature
EExNAS [23] is a first attempt to optimize an EENN with NAS. It accounts for only one EEC at a fixed position. Differently, ESAI [24] is a NAS for EENNs in a Cloud setting. The EENN is organized into two sub-parts, i.e., a lighter one is processed locally and another heavier one on cloud. The early exit mechanism is used to decide where to send the samples to units (e.g., local or remote server) of a technological ecosystem. HADAS [21] presents a framework to design the backbone network and EECs by a two-level optimization with two different genetic algorithms. On the first level, HADAS selects the optimal backbone neural networks by sampling from a Once-For-All (OFA) supernet [26] and taking into consideration the hardware constraints, while on the second level, it enhances the chosen backbones with the EECs by automating the choice of their number and placement, trains the obtained EENNs, and then selects the optimal EENNs. The EENNs are trained by using the Knowledge Distillation (KD) approach.
@@
A different approach suggests that the OFA can be enhanced and trained with early exits [27]. The work presents OFAv2, a novel OFA extending the OFAMobileNetv3 architecture by adding parallel blocks, additional dense skip connections, and implementing early exits. To support the redesign of the OFA, the training algorithm of the OFA, named Progressive Shrinking, has been expanded by adding two additional steps to handle the parallel blocks and the early exits, respectively. The NAS approach used in the experimental campaign relying on the novel OFA is evaluated on the Tiny Imagenet Dataset, a smaller version of Imagenet, which provides a significant increase in accuracy with respect to the original OFA architecture and maintains the computational efficiency.
@@
Currently, the NAS solutions for EENNs are very limited in jointly designing the backbones and the EECs and don’t take into account their computational demand. EDANAS [25] is the first framework for the automatic design of both the backbone architecture and the parameters managing the early exit mechanism of the EENNs so as to optimize both the accuracy of the classification tasks and the computational demand. The problem addressed by EDANAS is formulated as a joint optimization problem consisting in selecting the architecture of an EENN aiming at providing a high accuracy on a given task and reducing the MACs associated with the EE by using the genetic algorithm NSGA-II [28].
@@ III-A Early Exit Neural Networks
where $\circ$ denotes function composition. This is a generic definition which is agnostic about the internal topology of a block, and that applies to most neural networks. Moreover, we can apply a classification layer to each block of the model, which leads to multiple prediction functions:
@@
$L_{\text{joint}}(x,y)=l(f_{b}(x),y)+\sum_{i=1}^{b-1}\lambda_{e}\cdot\underbrace{\left[l(f_{i}(x),y)\right]}_{\text{$i$-th branch loss}}\,,$ (3)
where $l(\cdot,\cdot)$ is the loss (e.g. cross-entropy loss). Other training approaches are Layer-wise Training (LT), aiming at separately optimizing each layer in the NN architecture [29], and Knowledge Distillation (KD) where the backbone network is trained separately from the EECs and then the EECs are trained on top of it by relying on KD [30].
After training the model using the aforementioned approach, we can use this formulation to select the earliest possible exit which is assumed to output the correct prediction. This leads to many benefits such as the reduction of the inference time [1] [9], overfitting [3] and overthinking [1]. To do this we can use different early exit mechanisms, such as entropy, proposed in [1], the largest value of the softmax, or the so-called Score Margin (SM) [15] [31], which is the difference between the largest and the second largest value on the softmax on the prediction. These mechanisms are simple to implement, but they introduce a potential mismatch between training and test behaviours, since the early exit mechanism is not optimized on the training data.
@@
We divide the loss into two parts to avoid the intermediate branches, the second term, overshadowing the final output, the first term, which can be small if the model is highly confident in the early stage of the computation. This formulation allows us to halt the inference process at a given exit $i$ when the following criterion is met:
@@
$\mathcal{R}_{i}(x,y)=$ $-c_{i}(x)\log(\mathds{1}(y,\overline{f}_{i}(x))$ (9)
@@ III-B Neural Architecture Search
The second dimension is the Search Strategy, which establishes how to explore the search space. Common strategies are based on reinforcement learning (RL) [35, 36], gradient optimization [37], and evolutionary algorithms (EAs) [38, 39]. The latter, in particular the genetic methods, represents the most popular choice [40]. A typical genetic algorithm deployed in many NAS solutions is NSGA-II [28], which generates offsprings by using a specific type of crossover and mutation and then selects the next generation showing the best fitness according to nondominated-sorting and crowding distance comparison. Moreover, it is a multi-objective algorithm, i.e., being able to optimize jointly different figures of merit, and it is able to effectively handle constraints [41]. The constraints are handled by redefining the objective function [42], and the simplest scheme works as follows:
@@
The third dimension is the Performance Estimation Strategy, which refers to the process of estimating the performance to be optimized. The purpose of this family of methods is to avoid training every single neural network architecture in the search space which would be prohibitive in terms of computational time. Weight sharing is a popular estimation strategy, and in particular, Once-For-All (OFA) [26] represents a state-of-the-art. OFA trains a supernet comprehensive of many network configurations once and evaluates the candidate networks by fine-tuning from the supernet weights during the search of the NAS. The training of the OFA supernet is performed by a technique called progressive shrinking, which first trains the largest neural network with maximum depth, width, and kernel size, then progressively fine-tunes the OFA network to support smaller sub-networks that share weights with the larger ones. Another solution is the use of surrogate models such as a Gaussian Process [43] and Radial Basis Function [44]. For instance, MSuNAS [45] proposes a method called adaptive-switching which selects the best accuracy predictor among four possible surrogate models according to a correlation metric named Kendall’s Tau [46] at each iteration of the NAS process. Moreover, another technique called weight inheritance (network morphism) [47] evaluates the performance of a network obtained by an identity transformation (morphism) on a parent model by inheriting the weights of the parent model, i.e., being initialized with the same weights.
@@ IV The proposed NACHOS
This section introduces and details the presented NACHOS. More specifically, Section IV-A provides the problem formulation of NACHOS, Section IV-B provides an overall description of NACHOS, Section IV-C details the automatic placement of the EECs along the backbone, Section IV-D describes briefly the training phase by detailing the introduced regularization terms and the inference phase of the EENNs. Finally, Section IV-E introduces a novel figure of merit handling for the number of MACs of the EENN and the corresponding constraint enforced.
@@ IV-B The overall view of NACHOS
The fundamental part of NACHOS is the Search Strategy module. It adopts the genetic algorithm NSGA-II [28] to solve the bi-objective problem introduced in Equation (13), by jointly optimizing the objectives $\mathcal{S}_{A}(\widetilde{x})$ and $\mathcal{S}_{CM}(\widetilde{x})$ on the dataset $\mathcal{DS}$, and by evaluating network architectures candidates from $\Omega_{\widetilde{x}}$. The search process is iterative: at each iteration, the two surrogates model, computing $\mathcal{S}_{A}(\widetilde{x})$ and $\mathcal{S}_{M}(\widetilde{x})$, respectively, are chosen employing a mechanism called adaptive-switching, the same method used by MSuNAS [45], which selects the best surrogate model according to a correlation metric (i.e., Kendall’s Tau [46]). The surrogate models are trained by using the archive as the dataset. Then, a ranking of the candidate networks, based on $\mathcal{S}_{A}(\widetilde{x})$ and $\mathcal{S}_{CM}(\widetilde{x})$, is computed and a new set of candidates is obtained by NSGA-II and forwarded to the Trainer. The Trainer updates the archive, which becomes available for evaluation in the next iteration.
@@ IV-C The OFA Adapter module: automatic placement of EECs
The placement of the EECs is given by sampling from $\Omega_{\theta}$ a binary encoding configuration $\theta$ of length $\overline{B}-1$ being $\overline{B}$ a user-defined parameter representing the maximum number of EECs, distributed along the backbone equidistantly and including the final classifier. For instance, with $\overline{B}=5$, a sample $\theta=[1,0,1,0]$ extracted from $\Omega_{\theta}$ means that EECs are placed at $1/5$ and $3/5$ of the backbone.
The steps of the placement of the EECs along the backbone are shown in Algorithm 1.
@@ 12 Add $\widetilde{x}$ to $\Omega_{\widetilde{x}}$;
13 return $\Omega_{\widetilde{x}}$ ;
Algorithm 1 OFA Adapter update
Figure 2: The update of the $\Omega_{\widetilde{x}}$ performed by the OFA adapter. The EECs are placed along the backbone $x$ according to the configuration $\theta$ (Line 2). The vector of computational costs $\gamma$ is computed on $x$ and EECs (Line 3). In the For statement each EEC is scanned and the condition $\gamma_{i}\leq\gamma_{i+1}$ is checked (Lines 5-9). The last exit is the only one fixed a priori (i.e., basic design by default). If an EEC violates this condition, a Max Pooling layer is added through the operator AddMaxPooling()($\cdot$)( ⋅ ), where the window of the Max Pooling layer is the minimum s.t. this EEC satisfies the condition. Finally, the designed EECs are attached to the backbone leading to the creation of an EENN (Line 10) and the EENN is added to $\Omega_{\widetilde{x}}$ (Line 11).
@@ 8       Compute the average accuracies per class on the Support Set;
12Set $\widetilde{x}$ with the weights for the best validation loss;
@@
Figure 3: Firstly, the Trainer extracts the training and validation set from the dataset $DS$ (Line 1). Then, it performs fine-tuning on the backbone of the EENN for a number of epochs $\mu_{1}$ starting from the pretrained weights of the OFA supernet (Lines 2-3). Once trained the backbone, the Trainer trains the EENN with the soft exit mechanism presented in Section  III-A with two regularization terms, $L_{cost}$ and $L_{peak}$. In particular, we can distinguish between a phase where only a regularization $L_{cost}$ on computational costs is applied for a $\mu_{2}$ number of epochs (Lines 4-5) and a second phase where an additional regularization $L_{peak}$ for enhancing the accuracy is applied lasting $\mu_{3}$ number of epochs. This is done to trigger the latter regularization only after the EECs have been trained for a sufficient number of epochs. (Lines 6-10). We then select the weights corresponding to the best loss computed on the validation set (Line 11).
The training performed by the Trainer module is shown in Algorithm 2. The overall train loss function is the following:
@@
First, we examine the regularization to enforce the constraint on the complexity term. As introduced in EDANAS [25], the metric $F_{M}(\widetilde{x})$ can be computed as the sum of $\gamma_{i}$ weighted by the utilization of the corresponding $EEC_{i}$ and can be computed as follows:
@@
$\tilde{\gamma}_{i}(\widetilde{x})=c_{i}\gamma_{i}(\widetilde{x})+(1-c_{i}\tilde{\gamma}_{i+1})(\widetilde{x})$ (17)
@@
Second, we explain the regularization introduced to improve the accuracy forcing the confidences to follow the operating point of the EENN. First, part of the original training set, containing an equal number of samples for each class, is extracted. We call this extracted set as Support Set and the number of classes as $C$. The Support Set is used to compute a matrix $C\times B$ named Support Matrix. The Support Matrix contains the average accuracies per class computed for each epoch on top of the support set. For each class, the average accuracy is the ratio of the corrected predicted samples in the set of samples of a given class in the support set. Once defined the Support Matrix, we use a regression loss as a regularization term to force the local certainties (calculated on the batch samples) to follow the global accuracies (calculated on the Support Set). Basically, we compute the Mean Squared Error (MSE) on two vectors of $B$ values computed per each sample as follows:

### v2

@@ I Introduction
Neural Architecture Search (NAS) is a prominent research area [17, 18, 19] aiming at automating the design of a neural network by exploring different architectural configurations given a neural network topology, a task to be accomplished, and a dataset used for training and validation so as to provide the optimal architecture. The majority of works in this field focus on the design of deep neural network architectures while taking into account a specific target hardware. This family of NAS solutions is called Hardware-Aware NAS [20]. An emerging direction of these NAS solutions regards the use of constraints in the NAS exploration [21, 22, 23, 24]. Specifically, the works belonging to this research direction are able to consider both functional constraints (i.e., the type of operations that can be carried out in the neural network) and technological constraints (i.e., constraints on the computational and memory demand of the designed neural network in the exploration of the neural networks). The constraints are usually handled by a genetic algorithm applying a penalization to the fitness of the architectures not satisfying the hardware constraints. Interestingly, there are very few examples of NAS for EENNs in the current literature and these works are not able to provide an effective joint design of the backbone networks and the EECs [25, 26, 23, 27]. These NAS solutions are commented in Section II.
@@
  • a novel NAS solution able to co-design the backbone and the EECs in EENNs satisfying user-defined constraints by automatically placing the EECs in the backbone of an EENN accounting for their computational complexity;
@@
The rest of this paper is organized as follows. Section II introduces the related works regarding EENNs designed by NAS solutions. Section III provides the background. Section IV introduces the NACHOS framework developed in this study, while in Section V the experimental results aiming at evaluating the effectiveness of NACHOS are shown. Conclusions are finally drawn in Section VI. To facilitate comparisons and reproducibility, the source code of NACHOS is released to the scientific community as a public repository. [^1] https://github.com/AI-Tech-Research-Lab/CNAS
@@ II Related literature
EExNAS [25] is a first attempt to optimize an EENN with NAS. It accounts for only one EEC at a fixed position. Differently, ESAI [26] is a NAS for EENNs in a Cloud setting. The EENN is organized into two sub-parts, i.e., a lighter one is processed locally and another heavier one on cloud. The early exit mechanism is used to decide where to send the samples to units (e.g., local or remote server) of a technological ecosystem. Both works operate on medical applications, the first focusing on Myocardial Infarction and Human Activity Recognition, and the second on the classification of dermoscopic images. HADAS [23] presents a framework to design the backbone network, the EECs, and the dynamic voltage and frequency scaling settings, by a two-level optimization with two different genetic algorithms. On the first level, HADAS selects the optimal backbone neural networks by sampling from a Once-For-All (OFA) supernet [28] and taking into consideration the hardware constraints, while on the second level, it enhances the chosen backbones with the EECs by automating the choice of their number and placement, trains the obtained EENNs, and then selects the optimal EENNs. The EENNs are trained by using the Knowledge Distillation (KD) approach. This NAS takes 2-3 GPU days on a cluster of 32 GPUs on CIFAR-100.
@@
A different approach suggests that the OFA can be enhanced and trained with early exits [29]. The work presents OFAv2, a novel OFA extending the OFAMobileNetv3 architecture by adding parallel blocks, additional dense skip connections, and implementing early exits. To support the redesign of the OFA, the training algorithm of the OFA, named Progressive Shrinking, has been expanded by adding two additional steps to handle the parallel blocks and the early exits, respectively. The NAS approach used in the experimental campaign relying on the novel OFA is evaluated on the Tiny Imagenet Dataset, a smaller version of Imagenet, which provides a significant increase in accuracy with respect to the original OFA architecture and maintains the computational efficiency. All in all, OFAV2 performs the search on a previously trained supernet specifically designed with multi-exits.
@@
Currently, there are only few NAS solutions that jointly design the backbones and the EECs while taking into account the computational demand. EDANAS [27] is a framework for the automatic design of both the backbone architecture and the parameters managing the early exit mechanism of the EENNs so as to optimize both the accuracy of the classification tasks and the computational demand. The problem addressed by EDANAS is formulated as a joint optimization problem consisting in selecting the architecture of an EENN aiming at providing a high accuracy on a given task and reducing the MACs associated with the EE by using the genetic algorithm NSGA-II [30]. Similarly, HarvNet [31] addresses the joint design of backbone and EECs while meeting memory and energy constraints in dynamic harvesting environments. It relies on reinforcement learning to adjust the confidence threshold of each exit to maximize accuracy by considering the device’s runtime factors. This NAS takes 47 GPU days on CIFAR-10, since they integrate the runtime aspects of dynamic input mappings and dynamic harvesting environments in the design of NNs respectively.
Lastly, we emphasize that our work shows a hardware-aware NAS for EENNs on computer vision applications working on top of a standard single-exit OFA supernet independent of input mappings and the environment with a search time of 2 GPU days on a single GPU (NVIDIA A40). This overcomes some of the limitations of the previous works such as limited application scenarios, specific design of the supernet, and high computational requirements.
@@ III-A Early Exit Neural Networks
where $\circ$ denotes function composition. This is a generic definition which is agnostic about the internal topology of a block, and that applies to most neural networks. Moreover, we can apply a classification layer to each block i-th of the model, which leads to multiple prediction functions:
@@
$L_{\text{joint}}(x,y)=l(f_{b}(x),y)+\sum_{i=1}^{b-1}\lambda_{i}\cdot\underbrace{\left[l(f_{i}(x),y)\right]}_{\text{$i$-th branch loss}}\,,$ (3)
where $l(\cdot,\cdot)$ is the loss (e.g. cross-entropy loss), and $\lambda_{i}$ the weight associated to the i-th exit’s loss. Other training approaches are Layer-wise Training (LT), aiming at separately optimizing each layer in the NN architecture [32], and Knowledge Distillation (KD) where the backbone network is trained separately from the EECs and then the EECs are trained on top of it by relying on KD [33].
After training the model using the aforementioned approach, we can take advantage of the formulation proposed in Eq. 2 to select the earliest feasible exit which is assumed to output the correct prediction. This leads to many benefits such as the reduction of the inference time [1] [9], overfitting [3] and overthinking [1]. To do this we can use different early exit mechanisms, such as entropy, proposed in [1], the largest value of the softmax, or the so-called Score Margin (SM) [15] [34], which is the difference between the largest and the second largest value on the softmax on the prediction. These mechanisms are simple to implement, but they introduce a potential mismatch between training and test behaviours, since the early exit mechanism is not optimized on the training data.
@@
We divide the loss into two parts to avoid the intermediate branches, the second term, overshadowing the final output, the first term, which can be small if the model is highly confident in the early stage of the computation. This training approach allows us to halt the inference process at a given exit $i$ when the following criterion is met:
@@
$\mathcal{R}_{i}(x,y)=$ $-c_{i}(x)\log(\mathds{1}(y,\overline{f}_{i}(x))$
@@ III-B Neural Architecture Search
The second dimension is the Search Strategy, which establishes how to explore the search space. Common strategies are based on reinforcement learning (RL) [38, 39], gradient optimization [40], and evolutionary algorithms (EAs) [41, 42]. The latter, in particular the genetic methods, represents the most popular choice [43]. The Non-dominated Sorting Genetic Algorithm II (NSGA-II) is widely used in NAS solutions thanks to its ability to efficiently explore the search space [30]. It generates offspring through crossover and mutation operations, and then applies a non-dominated sorting procedure combined with a crowding distance mechanism to select the most optimal individuals for the next generation. This approach ensures that the selected architectures are not only characterized by high fitness but also maintain a diverse set of solutions, helping to avoid premature convergence. Moreover, it is a multi-objective algorithm, i.e., being able to optimize jointly different figures of merit, and it is able to effectively handle constraints [44]. The constraints are handled by redefining the objective function [45], and the simplest scheme works as follows:
@@
The third dimension is the Performance Estimation Strategy, which refers to the process of estimating the performance to be optimized. The purpose of this family of methods is to avoid training every single neural network architecture in the search space which would be prohibitive in terms of computational time. Weight sharing is a popular estimation strategy, and in particular, Once-For-All (OFA) [28] represents a state-of-the-art. OFA trains a supernet comprehensive of many network configurations once and evaluates the candidate networks by fine-tuning from the supernet weights during the search of the NAS. The training of the OFA supernet is performed by a technique called progressive shrinking, which first trains the largest neural network with maximum depth, width, and kernel size, then progressively fine-tunes the OFA network to support smaller sub-networks that share weights with the larger ones. While OFA provides three variants of CNN-based supernets, inspired by MobileNetV3, ProxylessNAS, and Resnet50 respectively, NASViT proposed a Vision Transformer-based supernet [46]. The choice of the supernet accounts for the different desired accuracy-efficiency trade-off in a target scenario. Another solution is the use of surrogate models such as a Gaussian Process [47] and Radial Basis Function [48]. For instance, MSuNAS [49] proposes a method called adaptive-switching which selects the best accuracy predictor among four possible surrogate models according to a correlation metric named Kendall’s Tau [50] at each iteration of the NAS process. Moreover, another technique called weight inheritance (network morphism) [51] evaluates the performance of a network obtained by an identity transformation (morphism) on a parent model by inheriting the weights of the parent model, i.e., being initialized with the same weights.
@@ IV The proposed NACHOS
This section introduces and details the presented NACHOS. More specifically, Section IV-A provides the problem formulation of NACHOS, Section IV-B provides an overall description of NACHOS, Section IV-C details the automatic placement of the EECs along the backbone, Section IV-D describes briefly the training phase by detailing the introduced regularization terms and the inference phase of the EENNs. Finally, Section IV-E introduces a novel figure of merit handling for the number of MACs of the EENN and the corresponding constraint enforced. To improve the readability, Table I contains all symbols we use to define the problem and our proposal.
@@ IV-B The overall view of NACHOS
Figure 2: The process of creating an EENN and how such models are used to populate the searching space. The best model is searched within this space. For further details, please refer to Section IV-C.
@@
The fundamental part of NACHOS is the Search Strategy module. It adopts the genetic algorithm NSGA-II [30] to solve the bi-objective problem introduced in Equation (12), by jointly optimizing the objectives $\mathcal{S}_{A}(\widetilde{x})$ and $\mathcal{S}_{CM}(\widetilde{x})$ on the dataset $\mathcal{DS}$, and by evaluating network architectures candidates from $\Omega_{\widetilde{x}}$. The search process is iterative: at each iteration, the two surrogates model, computing $\mathcal{S}_{A}(\widetilde{x})$ and $\mathcal{S}_{M}(\widetilde{x})$, respectively, are chosen employing a mechanism called adaptive-switching, the same method used by MSuNAS [49], which selects the best surrogate model according to a correlation metric (i.e., Kendall’s Tau [50]). The surrogate models are trained by using the archive as the dataset. Then, a ranking of the candidate networks, based on $\mathcal{S}_{A}(\widetilde{x})$ and $\mathcal{S}_{CM}(\widetilde{x})$, is computed and a new set of candidates is obtained by NSGA-II and forwarded to the Trainer. The Trainer updates the archive, which becomes available for evaluation in the next iteration. The search ends if a number of iterations $N_{iter}$ is reached or the hypervolume of the Pareto front is not changing for $N_{tol}$ iterations (convergence).
@@ IV-C The OFA Adapter module: automatic placement of EECs
TABLE I: List of the main symbols used in the experiments
Symbol Meaning
$\Omega_{x}$ Set of OFA backbones.
$\Omega_{\theta}$ Set of EEC hyperparameters.
$x$ Neural network architecture (backbone) sampled by $\Omega_{x}$.
$\theta$ The configuration of hyperparameters sampled by the EEC set.
$\epsilon$ The set of confidence thresholds.
$\widetilde{x}$ EENN sampled by $\Omega_{\widetilde{x}}$.
$F_{A}(\widetilde{x})$ The top1 accuracy of $\widetilde{x}$.
$F_{M}(\widetilde{x})$ The number of MACs of $\widetilde{x}$.
@@
The placement of the EECs is given by sampling from $\Omega_{\theta}$ a binary encoding configuration $\theta$ of length $\overline{B}-1$ being $\overline{B}$ a user-defined parameter representing the maximum number of EECs, distributed along the backbone equidistantly and including the final classifier. The uniform placement ensures that the computational complexity is distributed uniformly along the backbone since the layers have similar computational complexity. For instance, with $\overline{B}=5$, a sample $\theta=[1,0,1,0]$ extracted from $\Omega_{\theta}$ means that EECs are placed at $1/5$ and $3/5$ of the backbone.
The steps of the placement of the EECs along the backbone are shown in Algorithm 1, while the process is visually depicted in Figure 2. In this Algorithm, the EECs are placed along the backbone $x$ according to the configuration $\theta$ (Line 2). The vector of computational costs $\gamma$ is computed on $x$ and EECs (Line 3). In the For statement each EEC is scanned and the condition $\gamma_{i}\leq\gamma_{i+1}$ is checked (Lines 5-9). The last exit is the only one fixed a priori (i.e., basic design by default). If an EEC violates this condition, a Max Pooling layer is added through the operator AddMaxPooling()($\cdot$)( ⋅ ), where the window of the Max Pooling layer is the minimum s.t. this EEC satisfies the condition. Finally, the designed EECs are attached to the backbone leading to the creation of an EENN (Line 10) and the EENN is added to $\Omega_{\widetilde{x}}$ (Line 11).
@@ 12 Add $\widetilde{x}$ to $\Omega_{\widetilde{x}}$;
13
14return $\Omega_{\widetilde{x}}$ ;
Algorithm 1 OFA Adapter update: the update of the $\Omega_{\widetilde{x}}$ performed by the OFA adapter
@@ 9       Compute the average accuracies per class on the Support Set;
13Set $\widetilde{x}$ with the weights for the best validation loss;
@@
The training performed by the Trainer module is shown in Algorithm 2. Firstly, the Trainer extracts the training and validation set from the dataset $DS$ (Line 1). Then, it performs fine-tuning on the backbone of the EENN for a number of epochs $\mu_{1}$ starting from the pre-trained weights of the OFA supernet (Lines 2-3). Once trained the backbone, the Trainer trains the EENN with the soft exit mechanism presented in Section  III-A with two regularization terms, $L_{cost}$ and $L_{peak}$. In particular, we can distinguish between a phase where only a regularization $L_{cost}$ on computational costs is applied for a $\mu_{2}$ number of epochs (Lines 4-5) and a second phase where an additional regularization $L_{peak}$ for enhancing the accuracy is applied lasting $\mu_{3}$ number of epochs. This is done to trigger the latter regularization only after the EECs have been trained for a sufficient number of epochs. (Lines 6-10). We then select the weights corresponding to the best loss computed on the validation set (Line 11).
The overall train loss function is the following:
@@
First, we examine the regularization to enforce the constraint on the complexity term. We aim at training more the fraction of the EENN not violating the constraint. As introduced in EDANAS [27], the metric $F_{M}(\widetilde{x})$ can be computed as the sum of $\gamma_{i}$ weighted by the utilization of the corresponding $EEC_{i}$ and can be computed as follows:
@@
$\tilde{\gamma}_{i}(\widetilde{x})=c_{i}\gamma_{i}(\widetilde{x})+(1-c_{i})\tilde{\gamma}_{i+1}(\widetilde{x})$ (16)
@@
Second, we explain the regularization introduced to force the confidences to follow the operating point of the EENN. We aim at improving the accuracy of the EENN by tuning the confidences while considering the performance on a representative set. First, part of the original training set, containing an equal number of samples for each class, is extracted. We call this extracted set as Support Set and the number of classes as $C$. The Support Set is used to compute a matrix $C\times B$ named Support Matrix. The Support Matrix contains the average accuracies per class computed for each epoch on top of the support set. For each class, the average accuracy is the ratio of the corrected predicted samples in the set of samples of a given class in the support set. Once defined the Support Matrix, we use a regression loss as a regularization term to force the local certainties (calculated on the batch samples) to follow the global accuracies (calculated on the Support Set). Basically, we compute the Mean Squared Error (MSE) on two vectors of $B$ values computed per each sample as follows: